# On the Robustness of Fully-Spiking Neural Networks in Open-World Scenarios using Forward-Only Learning Algorithms

### v1

@@ Appendix A Reversal of Order in FF-SCP
This appendix formalizes the mathematical arguments that drive the reversal of the order in the scoring function. As pointed out in Subsection 3.2, the main limitation of the initial scoring function depicted in Equation (9) arises when, given a randomly selected sample from the ID dataset $X$, the following inequality holds:
$\min_{c\in C}d(\textbf{0},L_{c,c})\leq\mathbb{E}\left[\min_{c\in C}d(\mathcal{N}_{1}(X,c),L_{c,c})\right].$ (12)
@@
For instance, this reversal effect of the order predominantly occurs in OoD samples that were identified as negative inputs by the network, resulting in a near-zero latent state. Formally, let $\mathcal{N}$ be an accurate enough network, defined as a network in which the following expression almost always holds:
@@
where $\bm{X}_{\ominus}$ denotes a random negative sample and $\epsilon$ represents a sufficiently low value. If this property holds, then we can understand the geometric behavior of $\mathcal{L}_{c,p}$, where $p$ and $c$ represent non-equal classes from $\mathcal{C}$.
Lemma A.1.
The discrete set $\mathcal{L}_{c,p}$, where $c$ and $p$ are distinct elements from $\mathcal{C}$, has a near-zero mean norm after the initial filtering step of the FF-SCP algorithm in accurate enough networks.
Proof.
Initial elements from $\mathcal{L}_{c,p}$ are predominantly composed of elements with a norm less than $\epsilon$ due to Equation (13), and, consequently, have a norm close to 00, with only a few elements having larger norms. After the highly active filtering step of the FF-SCP, it will be ensured that almost always the latent space $\mathcal{L}_{c,p}$ will be composed of elements with near-zero norm, and, as such, the fact that $\mathcal{L}_{c,p}\sim\textbf{0}$ can be asserted. ∎
Given an accurate enough network, in which a negative sample $x_{\ominus}$ from an OoD distribution satisfies Equation (13), the scoring function can be simplified to:
$s(x_{\ominus})=\min_{p\in C}\sum_{c\in C}d(\mathcal{N}_{1}(x_{\ominus},p),\mathcal{L}_{c,p})\approx\min_{p\in C}\sum_{c\in C}d(\textbf{0},\mathcal{L}_{c,p}).$ (14)
Additionally, since $d(\textbf{0},\mathcal{L}_{c,p})\sim d(\textbf{0},\textbf{0})=0$ when $p\not=c$, the expression can be further simplified to:
$s(x_{\ominus})=\min_{p\in C}\sum_{c\in C}d(\textbf{0},\mathcal{L}_{c,p})\approx\min_{p\in C}d(\textbf{0},\mathcal{L}_{p,p}).$ (15)
This implies that for an OoD sample that remains with near-zero norms for all candidate embedded labels, the value of $s(x_{\ominus})$ will be approximately equal to the distance between the zero vector and the closest point among all the latent clusters $\mathcal{L}_{p,p}$, where $p\in\mathcal{C}$.
@@
$s(x_{\oplus})=\min_{p\in C}\sum_{c\in C}d(\mathcal{N}_{1}(x_{\oplus},p),\mathcal{L}_{c,p})\approx\min_{p\in C}d(\mathcal{N}_{1}(x_{\oplus},p),\mathcal{L}_{p,p}),$ (16)
@@
which implies that $s(x_{\oplus})$ is approximated by the minimal distance between the latent vector of $x_{\oplus}$ and the closest latent point in $L_{p}^{p}$. If the distance between points inside the cluster is sufficiently large, to the extent where it is surpasses the value $\min_{p\in C}d(\textbf{0},\mathcal{L}_{p,p})$, the value of $s(x_{\oplus})$ will be greater than $s(\ominus)$. The distance inside the latent clusters can become large due to several reasons, including the curse of dimensionality, insufficient sampling points for the latent clusters, and the tendency of FFA to push high activity points away from the zero vector, creating cluster spaces with a large diameter. While increasing the number of samples for the latent representation can enhance the distance accuracy, it comes at the expense of increased computational requirements.
Appendix B Latent State of FFA
This appendix offers additional insights into the structure of the latent activity of networks trained using the FFA and the enhanced representational capabilities it provides. To analyze this space, we rely on the trained models used for the OoD detection experiments, specifically the analog model using ReLU activation functions and the bounded goodness SNN. These two models have been selected as they encompass both activity regimes observed during the experimentation: the sparse latent space and the bounded latent space. To avoid redundancy in the latent space, and given the similar latent behavior arising from the different datasets, we employ random batches of data from the MNIST dataset to extract the latent spaces.
The plot showing the latent vectors obtained from the analog ReLU network is depicted in Figure 8. Consistent with the findings of Tosato et al. [4] and Yang [19], the latent vectors emerging from these networks are characterized by sparse activity regimes and high neural specialization. As observed in the images, most latent vectors are composed of only a small subset of active neurons, which represent key distinctive features of the different classes. For instance, given that the latent vectors of the images are ordered by class, it can be observed that the set of neurons that activate when presented with each input distribution is highly consistent among the batch. Similarly, each neuron is usually only active when presented with a small subset of the available input classes, often being active with only one.
Figure 8: Latent vector of the first layer of 256 random samples extracted from a ReLU activated network trained with FFA. Only the first 400 neurons are displayed. The samples are ordered by class.
Similarly, Figure 9 shows the latent activity from the network trained using FFA with the $G_{0}(\cdot)$ goodness function. In contrast to the previous network, the representations obtained showcase a denser activity regime, with latent vectors comprising a larger set of active neurons per instance. As in the previous case, the activity of the data follows class-based activation patterns, with each neuron specializing in a specific set of classes. However, in this instance, the set of active classes results in a broader selection, allowing for more classes to activate each neuron. Upon early experimentation, the emergence of these representations appears to be a byproduct of the bounded regimes of the goodness functions. Due to the inability to achieve unbounded goodness through a small set of unbound neurons, the network relies on a larger set of neurons for each sample. In contrast with the unbounded models, where the negative samples resulted in near-zero latent vectors, bounded models persist in the increased latent activity, resulting in negative latent vectors that contain a positive-like structure but within lower activity regimes. While the results in Table 1 indicate a small decrease in the accuracy of these networks, the increased bound of activity yields positive outcomes in terms of representability, as shown in the OoD detection results in Table 2.
Figure 9: Latent vector of the first layer of 256 random samples extracted from the SNN using $G_{0}(\cdot)$. Only the first 400 neurons are displayed. The samples are ordered by class.
To provide a more in-depth analysis of the geometrical properties of these latent vectors, we employed dimensionality reduction methods to present approximate depictions of the structure of the latent space. We relied on the T-SNE algorithm [42], using 1024 randomly chosen samples from MNIST, with latent vectors obtained by using all 10 available class labels. This approach allows us to explore the geometrical relationship between positive and negative samples, beyond the activity differences.
The latent space of the analog networks is depicted in Figure 10. Once again, our results align with those observed by other researchers: positive samples cluster together, farther from the zero vector, and are highly separated from negative samples. Similarly, the position of the negative set also appears to contain a clearly defined structure, with most samples staying close to the origin. It’s important to note that T-SNE is known to exhibit a partially distorted projection of the space, where distances do not necessarily represent exact mappings to the higher-dimensional original space. As such, our assertion on the closeness of negative samples to the origin of the space has been further analyzed manually by verifying their activity magnitude distribution. This analysis confirms that negative samples showcase lower latent activities, and instances farther from zero are still closer to zero than their positive counterparts.
Figure 10: T-SNE projection of the latent vectors from the first layer of the analog neural network. The latent space includes 1024 random samples from MNIST, incorporating both positive and negative labels. Positive latent vectors are illustrated using darker colors to differentiate them from negative samples.
Conversely, Figure 11 illustrates the space obtained by employing the spiking network with bounded goodness. One clear property that arises as a consequence of the denser latent activity is the higher clustering of the negative latent vectors into smaller subclusters. These subclusters correspond to samples obtained from pairs of real classes and embedded labels, showing how the algorithm discerns between different class selections based on their arguments. This latent geometry offers more precise results when measuring distances to OoD samples, as the clusters of class/label representatives provide closer distances among samples from the same distribution. This results in a more distinct separation between positive and negative samples in the space. However, this geometric property does not necessarily translate to enhanced accuracy, as the denser latent vectors and higher separation also reduce the specialization of neurons in the layer.
Figure 11: T-SNE projection of the latent vectors from the first layer of the spiking neural network. The latent space includes 1024 random samples from MNIST, incorporating both positive and negative labels. Positive latent vectors are illustrated using darker colors to differentiate them from negative samples.
Appendix C OoD Detection Results without Order Reversal Considerations
@@
This appendix presents the results obtained when employing the FF-SCP algorithm when employing the $s(\cdot)$ scoring function. As previously mention in Section 3.2, the function $s(\cdot)$ does not properly measure the distance in scenarios where the intra-cluster distance is greater than the distance between the zero vector and each sample of positive latent clusters, which is formally proven in Appendix A. As expected, the results from Table 3 evidence the reduced OoD detection accuracy when compared to the ones from Table 2, which relies on the refined scoring function $S(\cdot)$.
@@
Table 3: Results of the FF-SCP algorithm using the $s(\cdot)$ of our three proposed networks: the Analog FF-Network; the Unbounded FF-Network, denoted as $G_{\infty}(\cdot)$; and the Bounded FF-Network, denoted as $G_{0}(\cdot)$. Data from the SCP and ODIN has been extracted from the paper [24]. The ANN results are only shown as a baseline to compare ANN vs SNN.
@@
ID Dataset OOD Dataset FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN

### v2

@@ Appendix A Reversal of Order in FF-OoD
This appendix formalizes the mathematical arguments that drive the reversal of the order in the scoring function. As pointed out in Section 3.2, the main limitation of the initial scoring function depicted in Equation (9) arises when, given a randomly selected sample from the ID dataset $X$, the following inequality holds:
$\min_{c\in C}\widehat{d}(\textbf{0},L_{c,c})\leq\mathbb{E}\left[\min_{c\in C}\widehat{d}(\mathcal{N}_{1}(X,c),L_{c,c})\right].$ (12)
@@
For instance, this reversal effect of the order predominantly occurs in OoD samples that were identified as negative inputs by the network, resulting in a near-zero latent state. Formally, let $\mathcal{N}_{k}(\cdot,\cdot)$ be an low-normed negative network, defined as a network in which the following expression holds:
@@
where $\bm{X}_{\ominus}$ denotes a random negative sample and $\epsilon$ represents a sufficiently low value.
When networks hold this property, we can verify that the discrete set of latent representatives $\mathcal{L}_{c,p}$, given two distinct $c,p\in\mathcal{C}$, will have near-zero mean norm after the initial filtering step of the FF-OoD algorithm. This effect occurs as the original set of negative representatives, which is predominantly composed of elements with a norm less than $\epsilon$, will have all high-normed elements removed after the initial filtering step. Therefore, we can assert that $\mathcal{L}_{c,p}\sim\textbf{0}$ whenever $c\not=p$.
For notational purposes, let $\mathcal{N}_{1}(x_{\ominus},p)$ be denoted as $\bm{\ell}^{p}_{\ominus}$, and $\mathcal{N}_{1}(x_{\oplus},p)$ as $\bm{\ell}^{p}_{\oplus}$ Then, given a low-normed negative network, the scoring function can be simplified to:
$s(x_{\ominus})=\min_{p\in C}\sum_{c\in C}\widehat{d}(\bm{\ell}^{p}_{\ominus},\mathcal{L}_{c,p})\approx\min_{p\in C}\sum_{c\in C}\widehat{d}(\textbf{0},\mathcal{L}_{c,p}).$ (14)
Additionally, since $\widehat{d}(\textbf{0},\mathcal{L}_{c,p})\sim\widehat{d}(\textbf{0},\textbf{0})=0$ when $p\not=c$, the expression can be further simplified to:
$s(x_{\ominus})=\min_{p\in C}\sum_{c\in C}\widehat{d}(\textbf{0},\mathcal{L}_{c,p})\approx\min_{p\in C}\widehat{d}(\textbf{0},\mathcal{L}_{p,p}).$ (15)
This result implies that the score $s(x_{\ominus})$ of an OoD sample $x_{\ominus}$ with near-zero norm will be almost equal to the distance between the origin of the space to the closest point among the positive latent clusters $\mathcal{L}_{p,p}$, with $p\in\mathcal{C}$.
@@
$\min_{p\in C}\sum_{c\in C}\widehat{d}(\bm{\ell}^{p}_{\oplus},\mathcal{L}_{c,p})\approx\min_{p\in C}\widehat{d}(\bm{\ell}^{p}_{\oplus},\mathcal{L}_{p,p}),$ (16)
@@
which implies that $s(x_{\oplus})$ is approximated by the minimal distance between the latent vector of $x_{\oplus}$ and the closest latent point in $L_{p}^{p}$. If the distance between points inside the cluster is sufficiently large, to the extent where it is surpasses the value $\min_{p\in C}\widehat{d}(\textbf{0},\mathcal{L}_{p,p})$, the value of $s(x_{\oplus})$ will be greater than $s(\ominus)$, thereby showing the reversing effect in the order of the OoD scores.
Appendix B OoD Detection Results without Order Reversal Considerations
@@
This appendix presents the results obtained when employing the FF-OoD algorithm when employing the $s(\cdot)$ scoring function. As previously mention in Section 3.2, the function $s(\cdot)$ does not properly measure the distance in scenarios where the intra-cluster distance is greater than the distance between the zero vector and each sample of positive latent clusters, which is formally proven in A. As expected, the results from Table 4 evidence the reduced OoD detection accuracy when compared to the ones from Table 2, which relies on the refined scoring function $S(\cdot)$.
@@
Table 4: Results of the FF-OoD algorithm using the $s(\cdot)$ of our three proposed networks: the Analog FF-Network; the Unbounded FF-Network, denoted as $G_{\infty}(\cdot)$; and the Bounded FF-Network, denoted as $G_{0}(\cdot)$. Data from the SCP and ODIN has been extracted from the paper (Martinez-Seras et al., 2023). The ANN results are only shown as a baseline to compare ANN vs SNN.
@@
ID Dataset OOD Dataset FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN
@@
Appendix C Latent State of FFA
To analyze the latent space of models trained using FFA, we rely on the trained models used for the OoD detection experiments, specifically the analog model using ReLU activation functions and the bounded goodness SNN. These two models have been selected as they encompass both activity regimes observed during the experimentation: the sparse latent space and the bounded latent space. To avoid redundancy in the latent space, and given the similar latent behavior arising from the different datasets, we employ random batches of data from the MNIST dataset to extract the latent spaces.
The plot showing the latent vectors obtained from the analog ReLU network is depicted in Figure 5a. Consistent with the findings of Tosato et al. (2023) and Yang (2023), the latent vectors emerging from these networks are characterized by sparse activity regimes and high neural specialization. As observed in the images, most latent vectors are composed of only a small subset of active neurons, which represent key distinctive features of the different classes. For instance, given that the latent vectors of the images are ordered by class, it can be observed that the set of neurons that activate when presented with each input distribution is highly consistent among the batch. Similarly, each neuron is usually only active when presented with a small subset of the available input classes, often being active with only one.
Analogously, Figure 5b depicts the latent activity from the aforementioned SNN. This image showcases denser latent vectors, with a larger amount of active neurons per instance, still keeping the class-based activation patterns within the activity. Upon early experimentation, the emergence of these less sparse representations appears to be a byproduct of the bounded regimes of the goodness functions. Due to the inability to achieve large goodness values through a small set of unbound neurons, the network employs on a large set of neurons that ends up adding to higher goodness scores. Furthermore, bounded models create non-dormant negative samples, which highly contrast with the near-zero latent activity from negative samples in unbounded models. While the results in Table 1 indicate a small decrease in the accuracy from these networks, the increased bound of activity yields positive outcomes in terms of representability, as shown in the OoD detection results in Table 2.
(a) Analog Network.
(b) Spiking Network.
Figure 5: Latent vectors of the first layer of 256 random samples ordered by class extracted from networks trained with FFA. Only the first 250 neurons are displayed.
To provide a more in-depth analysis of the geometrical properties of these latent vectors, we employed dimensionality reduction methods to present approximate depictions of the structure of the latent space. We relied on the T-SNE algorithm (Van der Maaten and Hinton, 2008), using 1024 randomly chosen samples from MNIST, with latent vectors obtained by using all 10 available class labels. This approach allows us to explore the geometrical relationship between positive and negative samples, beyond the activity differences.
The latent space of the analog and the spiking networks are depicted in Figure 6a and Figure 6b, respectively. Once again, our results align with those observed by other researchers: positive samples cluster together, farther from the zero vector, and are highly separated from negative samples. Similarly, the position of the negative set also appears to contain a clearly defined structure, with most samples staying close to the origin. In the spiking case, one clear property that arises as a consequence of the denser latent activity is the clustering of negative latent vectors into additional subclusters, which are formed by samples belonging to pairs of real classes and embedded labels. This latent geometry offers more precise results when measuring distances to OoD samples, as the clusters of class/label representatives provide closer distances among samples from the same distribution. While this effect increases the separation between positive and negative samples, this geometric property does not necessarily translate to enhanced accuracy, as the denser latent vectors also reduces the specialization of neurons in the layer.
(a) Analog Network.
(b) Spiking Network.
Figure 6: T-SNE projection of the 10240 latent vectors extracted from the first layer of the analog neural network over the MNIST dataset. Positive vectors are illustrated using darker colors to differentiate them from negative samples.